# HodgeFormer: Transformers for Learnable Operators on Triangular Meshes through Data-Driven Hodge Matrices

### v4

@@ Abstract
Currently, prominent Transformer architectures applied on graphs and meshes for shape analysis tasks employ traditional attention layers that heavily utilize spectral features requiring costly eigenvalue decomposition-based methods. To encode the mesh structure, these methods derive positional embeddings, that heavily rely on eigenvalue decomposition based operations, e.g. on the Laplacian matrix, or on heat-kernel signatures, which are then concatenated to the input features. This paper proposes a novel approach inspired by the explicit construction of the Hodge Laplacian operator in Discrete Exterior Calculus as a product of discrete Hodge operators and exterior derivatives, i.e. $(L:=\star_{0}^{-1}d_{0}^{T}\star_{1}d_{0})$. We adjust the Transformer architecture in a novel deep learning layer that utilizes the multi-head attention mechanism to approximate Hodge matrices $\star_{0}$, $\star_{1}$ and $\star_{2}$ and learn families of discrete operators $L$ that act on mesh vertices, edges and faces. Our approach results in a computationally-efficient architecture that achieves comparable performance in mesh segmentation and classification tasks, through a direct learning framework, while eliminating the need for costly eigenvalue decomposition operations or complex preprocessing operations.
@@ 1 Introduction
Traditional Transformers-based approaches for 3D data typically follow a graph-based application paradigm. Vanilla Transformers are used, usually with node features as input, while structural information is incorporated through positional embeddings, usually via eigen decomposition of a Laplacian-based matrix [40, 23, 19].
@@
The rest of the paper is organized as follows: Sec. 2 presents the related work. Sec. 3 provides preliminaries on aspects relevant to the proposed approach. Sec. 4 presents the methodology, and 5 presents the experimental results. Finally, Sec. 6 discusses the outcomes of this work and concludes this paper.
@@ 2 Related Work
Spectral methods are based on the eigen-decomposition of discrete operators, typically the Laplace-Beltrami operator, utilizing the translation, rotation, and scale invariance of operations in the spectral domain. HodgeNet [36] introduces a learnable class of sparse operators on meshes, built from standard constructions in discrete exterior calculus. It employs standard geometry processing operations and Hodge star operators ($\star_{0}$ and $\star_{1}$) to encode per-vertex and per-edge features. Spectral approaches by [8] leverage the graph Laplacian for convolution, with subsequent works like CayleyNets [17] improving spectral filtering design. However, these methods struggle with computational complexity due to the high computational cost of eigen-decomposition. To avoid this limitation, Surface Networks [15] leverage Graph Neural Networks (GNNs) to learn polynomial expansions of both intrinsic (Laplacian) and extrinsic (Dirac) operators to capture mesh geometry, enabling them to process irregular mesh structures directly. Shape descriptors built on the Laplace-Beltrami operator have been particularly successful in capturing geometric properties. The Heat Kernel Signature (HKS) [38] captures multiscale geometric information through the heat diffusion process, providing an isometry-invariant point descriptor. The Wave Kernel Signature (WKS) [1] offers an alternative based on quantum mechanical principles, achieving better feature localization. Building on these descriptors, several neural architectures have emerged. DiffusionNet [35] leverages diffusion-based features to achieve robust segmentation and classification on partial and non-isometric shapes. Geodesic CNNs [26] introduce geometry-adapted convolutional and pooling layers that operate in the spectral domain, utilizing HKS and WKS features as input channels for learning intrinsic shape representations. LaplacianNet [31] introduces a deep learning framework for 3D meshes encoding mesh connectivity using Laplacian spectral analysis. The network uses Mesh Pooling Blocks (MPBs) to split the mesh surface into local pooling patches and aggregate both local and global features. A mesh hierarchy is built from fine to coarse using Laplacian spectral clustering, making the method robust to different triangulations and isometric transformations. The method is dependent on eigen-decomposition and spectral clustering.
@@
Transformer adaptations display potential in capturing global relationships in 3D meshes. The Laplacian Mesh Transformer [18] applies Transformer layers to Laplacian eigenvectors and geometric features. MeshMAE [21] introduces self-supervised learning through masked autoencoders for mesh understanding. MeshFormer [24] addresses scalability by combining mesh simplification with hierarchical graph-based attention mechanisms. Mesh Transformer (MeT) [40] proposes a two-stream architecture that processes triangle and cluster features in parallel, using Laplacian eigenvectors for positional encoding. Despite their effectiveness, these Transformer-based methods face significant computational challenges. The eigen-decomposition required for Laplacian features and the quadratic complexity of attention mechanisms limit their applicability to large-scale meshes, while the need for complex architectural components increases implementation and training difficulty.
@@ 3.1 Triangular meshes & Discrete Operators
In this work, we focus on triangular meshes $\mathcal{M(V,E,F)}$ with vertices $n_{v}=|\mathcal{V}|$, edges $n_{e}=|\mathcal{E}|$, and faces $n_{f}=|\mathcal{F}|$. In the context of Discrete Exterior Calculus (DEC) [9], a triangular mesh is seen as a 2-dimensional oriented simplicial complex. It is composed of oriented k-simplices for $k=\{0,1,2\}$ corresponding respectively to vertices, edges, and faces. On each set of k-simplices, we consider discrete differential k-forms $\Omega^{k}$, that is, functions that assign values to the elements of each set. In addition, we employ two important constructs:
@@
    The discrete exterior derivative $d_{k}$, along with its adjoint $d_{k}^{T}$, that is, the linear map $d_{k}:\Omega^{k}\rightarrow\Omega^{k+1}$ mapping discrete $k$-forms to discrete $k+1$ forms.
@@ 4.1 HodgeFormer Layer
We propose a novel deep learning architecture shown in Fig. 2, named “HodgeFormer”, that leverages the multi-head attention mechanism for learning data-driven Hodge Laplacian operators on k-forms based on the parametrization of the Hodge Star $\star_{k}$. Let $x_{k}\in\mathbb{R}^{n_{k}\times d_{k}}$ a k-form on mesh elements with $k\in\{v,e,f\}$, respectively, of size $n_{k}$ and let $W_{Q_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, $W_{K_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, and $W_{V_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$ the learnable projection matrices transforming $x_{k}$ to corresponding representations $Q_{k}$, $K_{k}$, and $V_{k}$ as in Eq. 4. For simplicity we set $d_{k}=d$ for $k\in\{v,e,f\}$, and $d_{h_{k}}=d_{h}=d/h$ for $h$ attention heads. Then we can parametrize the Hodge Star operator $\star_{k}$ as:
@@ 4.2 End-to-end Architecture
The input consists of features extracted from the mesh, along with the sparse oriented incidence matrices $d_{0}$ and $d_{1}$. Dedicated embedding layers, discussed in Sec. 4.5, operate separately on vertex, edge, and face features and map them to target latent dimensions $d_{v}$, $d_{e}$, and $d_{f}$ (Sec. 4.5). Then, the latent embeddings for each mesh element are transformed by a mixture of sequentially placed HodgeFormer and Transformer layers and updated accordingly. The final embeddings are passed to the task head, which maps them to the task-specific output dimensions. Putting together HodgeFormer layers with vanilla Transformer layers can be thought of as combining local operators with global mixers. Different mixing strategies of HodgeFormer and vanilla Transformer layers can be applied.
@@ 4.3 Sparse Attention
Let $x\in\mathbb{R}^{n\times d}$ a k-form on some mesh element, where $n$ the number of k-form elements and $d$ the feature dimension, and $Q,K,V$ the linearly mapped queries, keys and values respectively. For each k-element $i\in{1,\ldots,n}$, we define a sparsity pattern $S_{i}$, which is the set of key positions that the query at position $i$ can attend to. Then, the update of each k-form element feature vector $x_{i}$ through the sparse attention mechanism can be formalized as
@@ 4.4 Input Features and Positional Embeddings
HodgeFormer input tokens are directly formed from features on vertices, edges, and faces to enable the application of DEC-based formulations. In addition, positional information is directly incorporated in the architecture as xyz coordinates through the input features for each mesh element.
Specifically, we consider three types of input features: (i) coordinates, (ii) normals and (iii) areas: The vertex features $x_{v_{in}}$ consist of the point 3D coordinates, vertex normals, $n_{v_{i}}$ calculated as the weighted average of incident face normals $n_{f_{i}}$ and the vertex associated cell area, calculated as the weighted average of incident face areas. The edge features $x_{e_{in}}$ consist of the point 3D coordinates of edge vertices, and the vertices opposite to the respective edge, edge normals calculated as the average of incident vertex normals, and the lengths of the edges that belong to the incident faces. The face features $x_{f_{in}}$ consist of the point 3D coordinates of the face vertices ordered according to the face’s orientation, the face normal and the face area. These design choices aim to incorporate information associated both to the primal and the dual mesh. The effect of each type of input features is showcased in the ablation study in Tab. 5.
@@ 4.5 Embedding Layer
By performing a one-hop feature aggregation, we allow our network to take into account neighbors’ information before embedding the input features to the target dimension.
@@
LaplacianNet [31] mlp v Yes n/a n/a n/a 92.2% 94.2% 93,9%
Laplacian2Mesh [10] cnn v Yes 100.0% 91.5% 88.6% 94.6% 96.6% 95,0%
@@
Table 3: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[20] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of Milano et al. [27] with hard ground truth labels at faces, as well on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods. The abbreviation ”trns” denotes the transformer architecture. Notes: The EMNN (MC+H) and EGNN (MC+H) results shown in the table are as reported by [39].
clarifies that while EGNN and EMNN [39] theoretically do not need pre-calculated Heat Kernel Signatures (HKS), the authors report precomputation of HKS values for both the SHREC and Human Segmentation benchmark datasets
@@ 5.1 Evaluation Setup
HodgeFormer layers use sparse attention, where each input element attends over $\sqrt{n}$ elements, consisting of local neighbors and random connections in an $4:1$ ratio. Plain Transformer Layers follow a standard implementation except for the incorporation of a linear attention mechanism as in  [13]. All MLP layers use ReLU non-linearities where required. We adopt the pre-LN architecture variant with residual connections, placing Layer Normalization (LayerNorm) before the input to each layer block. Following [32], we remove scale and bias terms from LayerNorm. Dropout is employed as a regularization mechanism both in the HodgeFormer and the Transformer layers. Finally, in all experiments we employ the neighbor-aware embedding layer as described in Sec. 4.5.
@@ 5.3 Model Variations Study
In Tab. 4, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In Tab. 6, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. Let $s_{k}$ for $k\in{v,e,f}$ the number of neighbors each element of a k-form attends to. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In Tab. 5, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.
@@ 6 Discussion, limitations and conclusion
In this work, we draw connections between discrete Hodge operators and Transformer-style attention mechanisms and propose a general architecture for deep learning on triangular meshes named HodgeFormer. Our architecture omits the construction of commonly used techniques such as expensive spectral features, eigenvalue decomposition operations, and complex complementary structures. Despite these omissions, HodgeFormer retains its expressiveness while fully leveraging the hardware-friendly properties of the Transformer architecture. In addition, by employing sparse attention mechanisms, the HodgeFormer architecture results in an overall complexity of $O(n^{1.5}d)$. HodgeFormer was tested on classification and segmentation tasks, demonstrating performance comparable to state-of-the-art methods. Strategies such as patching in MeshMAE [21] or mesh simplification preprocessing in MeshWalker [16] are part of future research on large-scale meshes and large-scale unsupervised training. Compared to a standard Transformer, our architecture has higher memory requirements due to the use of multiple attention mechanisms per layer. This requirement is partially mitigated by the employment of sparse attention. Although the proposed architecture has low complexity and is fully parallelizable, its implementation relies heavily on gather operations. Investigation of preprocessing strategies for input meshes with reordering algorithms such as the Cuthill-McKee [4] may allow the use of contiguous operations and enable better hardware utilization.

### v5

@@ Abstract
Currently, prominent Transformer architectures applied to graphs and meshes for shape analysis tasks employ traditional attention layers that heavily utilize spectral features requiring costly eigenvalue decomposition-based methods. To encode the mesh structure, these methods derive positional embeddings that heavily rely on eigenvalue decomposition-based operations, e.g. on the Laplacian matrix, or on heat-kernel signatures, which are then concatenated to the input features. This paper proposes a novel approach inspired by the explicit construction of the Hodge Laplacian operator in Discrete Exterior Calculus as a product of discrete Hodge operators and exterior derivatives, i.e. $(L:=\star_{0}^{-1}d_{0}^{T}\star_{1}d_{0})$. We adjust the Transformer architecture in a novel deep learning layer that utilizes the multi-head attention mechanism to approximate Hodge matrices $\star_{0}$, $\star_{1}$, and $\star_{2}$ and learn families of discrete operators $L$ that act on mesh vertices, edges and faces. Our approach results in a computationally-efficient architecture that achieves comparable performance in mesh segmentation and classification tasks, through a direct learning framework, while eliminating the need for costly eigenvalue decomposition operations or complex preprocessing operations. Code and data are publicly available at https://github.com/hodgeformer/
@@ 1 Introduction
Traditional Transformer-based approaches for 3D data typically follow a graph-based application paradigm. Vanilla Transformers are used, usually with node features as input, while structural information is incorporated through positional embeddings, usually via eigenvalue decomposition of a Laplacian-based matrix [40, 23, 19].
@@
The rest of the paper is organized as follows: Sec. 2 presents the related work. Sec. 3 provides preliminaries on aspects relevant to the proposed approach. Sec. 4 presents the methodology, and Sec. 5 presents the experimental results. Finally, Sec. 6 discusses the outcomes of this work and concludes this paper.
@@ 2 Related Work
Spectral methods are based on the eigenvalue decomposition of discrete operators, typically the Laplace-Beltrami operator, utilizing the translation, rotation, and scale invariance of operations in the spectral domain. HodgeNet [36] introduces a learnable class of sparse operators on meshes, built from standard constructions in discrete exterior calculus. It employs standard geometry processing operations and Hodge star operators ($\star_{0}$ and $\star_{1}$) to encode per-vertex and per-edge features. Spectral approaches by [8] leverage the graph Laplacian for convolution, with subsequent works like CayleyNets [17] improving spectral filtering design. However, these methods struggle with computational complexity due to the high computational cost of eigenvalue decomposition. To avoid this limitation, Surface Networks [15] leverage Graph Neural Networks (GNNs) to learn polynomial expansions of both intrinsic (Laplacian) and extrinsic (Dirac) operators to capture mesh geometry, enabling them to process irregular mesh structures directly. Shape descriptors built on the Laplace-Beltrami operator have been particularly successful in capturing geometric properties. The Heat Kernel Signature (HKS) [38] captures multiscale geometric information through the heat diffusion process, providing an isometry-invariant point descriptor. The Wave Kernel Signature (WKS) [1] offers an alternative based on quantum mechanical principles, achieving better feature localization. Building on these descriptors, several neural architectures have emerged. DiffusionNet [35] leverages diffusion-based features to achieve robust segmentation and classification on partial and non-isometric shapes. Geodesic CNNs [26] introduce geometry-adapted convolutional and pooling layers that operate in the spectral domain, utilizing HKS and WKS features as input channels for learning intrinsic shape representations. LaplacianNet [31] introduces a deep learning framework for 3D meshes encoding mesh connectivity using Laplacian spectral analysis. The network uses Mesh Pooling Blocks (MPBs) to split the mesh surface into local pooling patches and aggregate both local and global features. A mesh hierarchy is built from fine to coarse using Laplacian spectral clustering, making the method robust to different triangulations and isometric transformations. The method depends on eigenvalue decomposition and spectral clustering.
@@
Transformer adaptations display potential in capturing global relationships in 3D meshes. The Laplacian Mesh Transformer [18] applies Transformer layers to Laplacian eigenvectors and geometric features. MeshMAE [21] introduces self-supervised learning through masked autoencoders for mesh understanding. MeshFormer [24] addresses scalability by combining mesh simplification with hierarchical graph-based attention mechanisms. Mesh Transformer (MeT) [40] proposes a two-stream architecture that processes triangle and cluster features in parallel, using Laplacian eigenvectors for positional encoding. Despite their effectiveness, these Transformer-based methods face significant computational challenges. The eigenvalue decomposition required for Laplacian features and the quadratic complexity of attention mechanisms limit their applicability to large-scale meshes, while the need for complex architectural components increases implementation and training difficulty.
@@ 3.1 Triangular meshes & Discrete Operators
In this work, we focus on triangular meshes $\mathcal{M(V,E,F)}$ with vertices $n_{v}=|\mathcal{V}|$, $n_{e}=|\mathcal{E}|$ edges, and $n_{f}=|\mathcal{F}|$ faces. In the context of Discrete Exterior Calculus (DEC) [9], a triangular mesh is seen as a 2-dimensional oriented simplicial complex. It is composed of oriented $k$-simplices for $k\in\{0,1,2\}$ corresponding respectively to vertices, edges, and faces. On each set of $k$-simplices, we consider discrete differential $k$-forms $\Omega^{k}$, that is, functions that assign values to the elements of each set. In addition, we employ two important constructs:
@@
    The discrete exterior derivative $d_{k}$ along with its adjoint $d_{k}^{T}$, that is, the linear map $d_{k}:\Omega^{k}\rightarrow\Omega^{k+1}$ mapping discrete $k$-forms to discrete $k+1$ forms.
@@ 4.1 HodgeFormer Layer
We propose a novel deep learning architecture shown in Fig. 2, named “HodgeFormer”, that leverages the multi-head attention mechanism for learning data-driven Hodge Laplacian operators on $k$-forms based on the parametrization of the Hodge Star $\star_{k}$. Let $x_{k}\in\mathbb{R}^{n_{k}\times d_{k}}$ be a $k$-form on mesh elements with $k\in\{v,e,f\}$, respectively, of size $n_{k}$ and let $W_{Q_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, $W_{K_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, and $W_{V_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$ the learnable projection matrices transforming $x_{k}$ to corresponding representations $Q_{k}$, $K_{k}$, and $V_{k}$ as in Eq. 4. For simplicity, we set $d_{k}=d$ for $k\in\{v,e,f\}$, and $d_{h_{k}}=d_{h}=d/h$ for $h$ attention heads. Then we can parametrize the Hodge Star operator $\star_{k}$ as:
@@ 4.2 End-to-end Architecture
The input consists of features extracted from the mesh, along with the sparse oriented incidence matrices $d_{0}$ and $d_{1}$. Dedicated embedding layers, discussed in Sec. 4.5, operate separately on vertex, edge, and face features and map them to target latent dimensions $d_{v}$, $d_{e}$, and $d_{f}$. Then, the latent embeddings for each mesh element are transformed by a mixture of sequentially placed HodgeFormer and Transformer layers and updated accordingly. The final embeddings are passed to the task head, which maps them to the task-specific output dimensions. Putting together HodgeFormer layers with vanilla Transformer layers can be thought of as combining local operators with global mixers. Different mixing strategies of HodgeFormer and vanilla Transformer layers can be applied.
@@ 4.3 Sparse Attention
Let $x\in\mathbb{R}^{n\times d}$ be a $k$-form on some mesh element, where $n$ is the number of $k$-form elements and $d$ is the feature dimension, and $Q,K,V$ are the linearly mapped queries, keys and values respectively. For each $k$-element $i\in\{1,\ldots,n\}$, we define a sparsity pattern $S_{i}$, which is the set of key positions that the query at position $i$ can attend to. Then, the update of each $k$-form element feature vector $x_{i}$ through the sparse attention mechanism can be formalized as
@@ 4.4 Input Features and Positional Embeddings
HodgeFormer input tokens are directly formed from features on vertices, edges, and faces to enable the application of DEC-based formulations. In addition, positional information is directly incorporated in the architecture as xyz coordinates through the input features for each mesh element. Specifically, we consider three types of input features: (i) coordinates, (ii) normals and (iii) areas: The vertex features $x_{v_{in}}$ consist of the point 3D coordinates, vertex normals, $n_{v_{i}}$ calculated as the weighted average of incident face normals $n_{f_{i}}$ and the vertex associated cell area, calculated as the weighted average of incident face areas. The edge features $x_{e_{in}}$ consist of the point 3D coordinates of edge vertices, and the vertices opposite to the respective edge, edge normals calculated as the average of incident vertex normals, and the lengths of the edges that belong to the incident faces. The face features $x_{f_{in}}$ consist of the point 3D coordinates of the face vertices ordered according to the face’s orientation, the face normal and the face area. These design choices aim to incorporate information associated both to the primal and the dual mesh. The effect of each type of input features is showcased in the ablation study in Tab. 5.
@@ 4.5 Embedding Layer
By performing one-hop feature aggregation, we allow our network to take into account neighbors’ information before embedding the input features to the target dimension.
@@
LaplacianNet [31] mlp v Yes n/a n/a n/a 92.2% 94.2% 93.9%
Laplacian2Mesh [10] cnn v Yes 100.0% 91.5% 88.6% 94.6% 96.6% 95.0%
@@
Table 3: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[20] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of Milano et al. [27] with hard ground truth labels at faces, as well as on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods. The abbreviation ”trns” denotes the transformer architecture. Notes: The EMNN (MC+H) and EGNN (MC+H) results shown in the table are as reported by [39].
clarifies that while EGNN and EMNN theoretically do not need pre-calculated Heat Kernel Signatures (HKS), the authors report precomputation of HKS values for both the SHREC and Human Segmentation benchmark datasets [39, Sec.5,p.5].
@@ 5.1 Evaluation Setup
HodgeFormer layers use sparse attention, where each input element attends over $\sqrt{n}$ elements, consisting of local neighbors and random connections in a $4:1$ ratio. Plain Transformer Layers follow a standard implementation except for the incorporation of a linear attention mechanism as in  [13]. All MLP layers use ReLU non-linearities where required. We adopt the pre-LN architecture variant with residual connections, placing Layer Normalization (LayerNorm) before the input to each layer block. Following [32], we remove scale and bias terms from LayerNorm. Dropout is employed as a regularization mechanism both in the HodgeFormer and the Transformer layers. Finally, in all experiments we employ the neighbor-aware embedding layer as described in Sec. 4.5.
@@ 5.3 Model Variations Study
In Tab. 4, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In Tab. 6, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. Let $s_{k}$ be the number of neighbors each element of a $k$-form attends to for $k\in\{v,e,f\}$. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In Tab. 5, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.
@@ 6 Discussion, limitations and conclusion
In this work, we draw connections between discrete Hodge operators and Transformer-style attention mechanisms and propose a general architecture for deep learning on triangular meshes named HodgeFormer. Our architecture omits the construction of commonly used techniques such as expensive spectral features, eigenvalue decomposition operations, and complex complementary structures.
Despite these omissions, HodgeFormer retains its expressiveness while fully leveraging the hardware-friendly properties of the Transformer architecture. In addition, by employing sparse attention mechanisms, the HodgeFormer architecture results in an overall complexity of $O(n^{1.5}d)$. HodgeFormer was tested on classification and segmentation tasks, demonstrating performance comparable to state-of-the-art methods. Strategies such as patching in MeshMAE [21] or mesh simplification preprocessing in MeshWalker [16] are part of future research on large-scale meshes and large-scale unsupervised training.
Compared to a standard Transformer, our architecture has higher memory requirements due to the use of multiple attention mechanisms per layer. This requirement is partially mitigated by the employment of sparse attention. Although the proposed architecture has low complexity and is fully parallelizable, its implementation relies heavily on gather operations. Investigation of preprocessing strategies for input meshes with reordering algorithms such as the Cuthill-McKee [4] may allow the use of contiguous operations enabling better hardware utilization.
Acknowledgements
This research was supported by TUM Innovation Network CoConstruct N2201 and TUM Georg Nemetschek Institute - Artificial Intelligence for the Built World.